OpenStudy (anonymous):

Sam is observing the velocity of a car at different times. After two hours, the velocity of the car is 54 km/h. After four hours, the velocity of the car is 58 km/h. Part A: Write an equation in two variables in the standard form that can be used to describe the velocity of the car at different times. Show your work and define the variables used. (5 points) Part B: How can you graph the equation obtained in Part A for the first six hours? (5 points) @ganeshie8 @perl

OpenStudy (anonymous):

v=50+2t is part A correct?

OpenStudy (anonymous):

were v is velocity and t is time

OpenStudy (agl202):

V=50+2t V is velocity ..... T is time in hours!!! Hope this helps!!! : )

OpenStudy (anonymous):

but what is it talking about in standard form perl?

OpenStudy (agl202):

Great job @knightmare6!!!

OpenStudy (anonymous):

and it says show your work

OpenStudy (perl):

I think y = mx + b is fine for standard form

OpenStudy (perl):

there is a special 'standard form' Ax + bY = C , but it doesn't seem appropriate here

OpenStudy (anonymous):

okay but how did i get 50 i really dont know how i did

OpenStudy (perl):

first find the slope

OpenStudy (anonymous):

i did that

OpenStudy (anonymous):

(2,54) and (4,58)

OpenStudy (perl):

you are given two points (time , velocity)

OpenStudy (perl):

right

OpenStudy (anonymous):

which is y2-y1/x2-x1 which is 2

OpenStudy (perl):

two ordered pairs , i should say

OpenStudy (anonymous):

but now how do i get 50

OpenStudy (perl):

right, so far you have y = 2t + b

OpenStudy (anonymous):

v=2t+b

OpenStudy (perl):

you plug in either point and solve for b, the y intercept

OpenStudy (perl):

or the 'v intercept'

OpenStudy (anonymous):

so how do i do that

OpenStudy (agl202):

You are given : After two hours, the velocity of the car is 54 km/h. After four hours, the velocity of the car is 58 km/h. that means (2, 54) and (4, 58) are two points on the required graph and u can use these to write the equation for velocity in slope-intercept form

OpenStudy (perl):

plug in (2,54) for x and y

OpenStudy (perl):

54 = 2(2) + b

OpenStudy (perl):

here x = t , y = v

OpenStudy (anonymous):

oh okay i understand

OpenStudy (perl):

:)

OpenStudy (anonymous):

okay so now for Part B;

OpenStudy (anonymous):

do i just plug 6 into t?

OpenStudy (perl):

for part b you need to graph

OpenStudy (perl):

i have just the program for that

OpenStudy (agl202):

Plug in 0 for t and then 6 for t...you get (0,50) and (6,62) Connect those 2 points and you have your graph for the first 6 hours

OpenStudy (agl202):

Hope this helps ya!!! : )

OpenStudy (perl):

https://www.desmos.com/calculator/zlsc4g9tsf you have zoom out (scroll wheel down) to see the graph

OpenStudy (anonymous):

@perl im confused

OpenStudy (anonymous):

do i just plug in 0 and 6 for t?

OpenStudy (perl):

you can make a table

OpenStudy (perl):

or plug in 0 and 6 , yes , thats two values of the table

OpenStudy (anonymous):

for 0-6?

OpenStudy (perl):

for a line, 2 points are enough

OpenStudy (perl):

but it doesnt hurt to plot more points , helps to draw it

OpenStudy (anonymous):

so for part B i can say you can plug in 0 and 6 for t

OpenStudy (anonymous):

Is that a yes or no

OpenStudy (perl):

yes

OpenStudy (perl):

Part B: `How can you graph` the equation obtained in Part A for the first six hours? (5 points)

OpenStudy (perl):

you can pick t = 0, t = 6